I really need a refresher here! Please help :)
a)solve the equation symbolically
b)classify the equation as a contradiction, an identity, or a conditional equation.
7 - 9z = 2(3 - 4z) - z

Answer by scott8148(6628)

distributing (the 2) ___ 7 - 9z = 6 - 8z - z

7 - 9z = 6 - 9z ___ there is obviously something amiss

this would be classified as a contradiction
___ in order for the equality to be true , z must have different (contradictory) values on either side of the equation